OpenStudy (anonymous):

if f is the function defined by f(x) = (x^2+4x)^(1/3) and g is the antiderivative of f such that g(5) = 7, then g(1) =?

OpenStudy (anonymous):

I'm tutoring an AP calc student, and this one sort of stumped me, honestly. The method that seemed to make sense to me would be something like: 1) take antiderivative of f. 2) Use g(5) = 7 to solve for the constant of integration 3) Use the particular solution to solve for g(1) However, I got stuck at taking the antiderivative of f(x). Am I missing some obvious way of taking that antiderivative? How do you think this problem is meant to be solved? This was a calculator problem, so my student seemed to think the he could use his ti89 to take the antiderivative, although we didn't have one with us, so I wasn't able to try that.

myininaya (myininaya):

\[g(x)=\int\limits _0^x (t^2+4t)^\frac{1}{3} dt +C \\ g(5)=\int\limits_0^5(t^2+4t)^\frac{1}{3}+C=7 \\ \text{ use a calculator to solve for the definite integral } \] So say we get some number N. \[g(5)=N+C=7 \\ N+C=7 \\ C=7-N\] whatever N is you got from the calculator \[g(1)=\int\limits_0^1(t^2+4t)^\frac{1}{3} dt+7-N\] use a calculator again for the definite integral

OpenStudy (anonymous):

That makes a lot of sense, myininaya. I wasn't aware the ti89 would do definite integrals like that. That's helpful.

myininaya (myininaya):

I assumed it could because ti-83 could and ti-89>ti-83. :p

myininaya (myininaya):

http://www.batesville.k12.in.us/physics/calcnet/calculator/ti_89/definite_integral.htm I think this explains how to use it if the student doesn't know how to use it.